Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB SQS SXP XO Migrated2Zulip Archived2GitlabPages Archived2PDF not2MigrateOrArchive MigratedByGroup
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip DET Sample Environment SPB_mig
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 168     Entry time: 30 Apr 2023, 21:59
Author: Raphael de Wijn 
Group: SPB 
Subject: Shift summary 

SPB/SFX In-house research 4978 shift 3 - end

X-ray delivery

  • 9.3 keV 352 pulses / train, relatively stable intensity ~1.9 mJ / pulse
  • beam down ~30 min

Optical laser delivery

  • None

Achievements/Observations

  • Collected several runs for the in-line crystallisation project but some difficulties to inject, probably a coating issue
  • Saw a lot of crystals

Issues

  • Realised that there seems to be a vacuum leak from the load lock valve. Optimal pressure is only achieved when the load lock pump is running.
ELOG V3.1.4-7c3fd00